OpenStudy (anonymous):

Match each pair of equivalent ratios: Term: 1. 10/25 2. 6to10 3. 13/20 4. 15:40 Definition: A. 3/4 B. 6to16 C. 2:5 D. 65% I really need somebodies help! Please and Thank you <3

OpenStudy (anonymous):

1-C,3-D,4-B

OpenStudy (anonymous):

and for 2, the option is wrong

OpenStudy (anonymous):

1---- 10/25 ------ both numerator and denominator has the greatest common factor as 5, so we devide both by 5 to get 2/5

OpenStudy (anonymous):

There is no 2/5 option only 3/4 :/

OpenStudy (anonymous):

13/20------ numerator and denominator do not have any factor common, so this implies it is the simplest form, so we can say that answer will be in the in percentage form, so 13/20= 13*100/20=65%

OpenStudy (anonymous):

whhat do u want, we have to match the 1 with one of the correct answer, and option C is 2/5 only :/

OpenStudy (anonymous):

15/40= 3/8=6/ 16

OpenStudy (anonymous):

I really don't know what to do so I'm just going with your answers lol

OpenStudy (anonymous):

rest assured, u r right in followwing me

OpenStudy (anonymous):

So, just to make sure I have this all correct, 1 is c, 2 is a, 3 is d, and 4 is b?

OpenStudy (anonymous):

ya but 2 is incorrct, the answer for 2 is not there in the option